Shift report 20030204 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Feb 04 2003 - 16:05:57 EST

  • Next message: Apache: "Shift report 20030204 16:00-24:00"
    ------------------------------------------------------------
    Date: Tue Feb  4 16:05:57 2003, Shift: 16:00-24:00
    Supervisor: ykl, with: 
    Log book#13 pages 147 - 147
    
    Shift Summary:
    ------------------------------------------------------------
    1. Good beam until 11:15. Magnet and control problem
    for the rest of the time.
    
    2.
    
    MRS 45 deg 500A, FS 8 degree 1/8B.
       Run 7779 (MRS 290K, FS 84K).
       Run 7780 (MRS 124K, FS 36K).
    
    Plan for:
       500K at MRS 40 deg 1050A,
       continue for 130K at FS 8 deg 1/8A.
    
    Ped and DVM
       Run 7781(non-TPC Ped run).
       Run 7782(very short TPC ped run).
       Run 7783(TPC ped run).
       Run 7784(DVM run).
    
    3. No detector problem.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Feb 04 2003 - 16:06:25 EST